What is the answer to 6/8-2/4
3/4-1/8
4/5-1/2
3/4-1/3
5/6-1/3
5/6-2/3
3/5-1/10
1/2-1/3
4/5-2/10
1/2-2/5
2/3-1/4
2/3-1/5

Found 2 solutions by DeepaEvangeline, JimMarshall:
Answer by DeepaEvangeline(15)

6/8-2/4 = (6-4)/8 = 2/8 = 1/4

3/4-1/8 = (6-1)/8 = 5/8

4/5-1/2 = (8-5)/10 = 3/10

3/4-1/3 = (9-4)/12 = 5/12

5/6-1/3 = (5-2)/6 = 3/6 = 1/2

5/6-2/3 = (5-4)/6 = 1/6

3/5-1/10 = (6-1)/10 = 5/10 = 1/2

1/2-1/3 = (3-2)/6 = 1/6

4/5-2/10 = 4/5-1/5 = 3/5

1/2-2/5 = (5-4)/10 = 1/10

2/3-1/4 = (8-3)/12 = 5/12

2/3-1/5 = (10-3)/15 = 7/15

Answer by JimMarshall(82)

6/8-2/4 = (6-4)/8 = 2/8 = 1/4
3/4-1/8 = (6-1)/8 = 5/8
4/5-1/2 = (8-5)/10 = 3/10
3/4-1/3 = (9-4)/12 = 5/12
5/6-1/3 = (5-2)/6 = 3/6 = 1/2
5/6-2/3 = (5-4)/6 = 1/6
3/5-1/10 = (6-1)/10 = 5/10 = 1/2
1/2-1/3 = (3-2)/6 = 1/6
4/5-2/10 = 4/5-1/5 = 3/5
1/2-2/5 = (5-4)/10 = 1/10
2/3-1/4 = (8-3)/12 = 5/12
2/3-1/5 = (10-3)/15 = 7/15